OpenStudy (anonymous):

what is the vertex form of y=-1(x-8)^2-59

OpenStudy (anonymous):

(8,-59)

OpenStudy (anonymous):

how did you get -59 lol i got -51

OpenStudy (whpalmer4):

@sourwing that's the vertex, not to be confused with vertex form... @prostudent vertex form is \(y = a(x-h)^2+k\) with the vertex at \((h,k)\)

OpenStudy (anonymous):

it's already in the vertex form btw

OpenStudy (anonymous):

yeah, i read too fast.

OpenStudy (whpalmer4):

standard form is what you get if you expand it out: \[y = ax^2+bx+c\]in this case \[y = -x^2+16x-123\]

OpenStudy (anonymous):

wait so its 8 and -51 for the vertex?

OpenStudy (whpalmer4):

No. You can read the vertex off directly if you have it in vertex form. How are you coming up with -51?

OpenStudy (whpalmer4):

Here's the pattern: \[y = a(x-h)^2+k\] Here's your formula: \[y=-1(x-8)^2-59\] it should be apparent that \[a=-1\]\[-h=-8\]\[k=-59\]

OpenStudy (whpalmer4):

so the vertex is at \((h,k) = (8,-59)\)

OpenStudy (anonymous):

o i meant y=-x^2-16x-59 i typed it in wrong :(

OpenStudy (whpalmer4):

So the formula is supposed to be \[y = -x^2-16x-59\]and you want vertex form? I'd complete the square after factoring out a -1. \[y = -1(x^2+16x + 59)\]\[y = -1(x^2+16x+64 - 5)\]Can you take it from here?